OpenStudy (anonymous):

2x-3 divided by 4x^4 - 8x^3 + 6x^2 - 3x + 2

OpenStudy (anonymous):

can you factor out the second part where you have something like 2x-3( w/e) so you can have \[\frac{2x-3}{2x-3(w/e)}= \frac{1}{(w/e)}\]

OpenStudy (anonymous):

is this possible?